A body of mass 10kg and initially at rest is subjected to force of 20N for 1s. Calculate the change in kinetic energy
1
Expert's answer
2019-07-05T09:11:48-0400

Let's first find the acceleration of the body from the Newton's Second Law of Motion:


"F = ma,""a = \dfrac{F}{m} = \dfrac{20 N}{10 kg} = 2 \dfrac{m}{s^2}."

Then, we can find the final velocity of the body from the kinematic equation:


"v_f = v_i + at,"

here, "v_f" is the final velocity of the body, "v_i = 0" is the initial velocity of the body, "a" is the acceleration of the body and "t" is time.

Then, we get:


"v_f = v_i + at = 0 \dfrac{m}{s} + 2 \dfrac{m}{s^2} \cdot 1 s = 2 \dfrac{m}{s}."

Finally, we can find the change in kinetic energy:


"\Delta KE = KE_f - KE_i,""\Delta KE = \dfrac{1}{2}mv_f^2 - \dfrac{1}{2}mv_i^2 = \dfrac{1}{2}m(v_f^2 - v_i^2),""\Delta KE = \dfrac{1}{2} \cdot 10kg \cdot ((2 \dfrac{m}{s})^2 - (0 \dfrac{m}{s})^2) = 20 J."

Answer:

"\Delta KE = 20 J."